OpenStudy (anonymous):

Variable y varies directly with variable x, and y = 4 when x = 7. What is the value of x when y = 12? i think it is 21

OpenStudy (alexandervonhumboldt2):

so if y=4 =>x=7 y=12 =>x=?

OpenStudy (anonymous):

? what do you mean

OpenStudy (alexandervonhumboldt2):

x/12=7/4

OpenStudy (alexandervonhumboldt2):

x=7/4*12=7*3=21

OpenStudy (anonymous):

yeah

OpenStudy (alexandervonhumboldt2):

so x=21 when y=12

OpenStudy (anonymous):

yes thank you

OpenStudy (alexandervonhumboldt2):

np

OpenStudy (mathstudent55):

Direct variation is expressed as: \(y = kx\) where k is a constant (a number). Start with the direct variation equation: \(y = kx\) Now plug in the given point \((7, 4)\) to find the value of k: \(4 = k \times 7\) Which gives \(k = \dfrac{4}{7}\) Now you plug in the value of k in the direct variation equation: \(y = \dfrac{4}{7} x\) The question now is what x value do you have when y = 12? \(12 = \dfrac{4}{7}x\) Multiply both sides by 7/4 to get: x = 21